# Can Robots “Taste” Grapes? Estimating SSC with Simple RGB Sensors

### v1

@@ Abstract
In table grape cultivation, harvesting depends on accurately assessing fruit quality. While some characteristics, like color, are visible, others, such as Soluble Solid Content (SSC), or sugar content measured in degrees Brix (°Brix), require specific tools. SSC is a key quality factor that correlates with ripeness, but lacks a direct causal relationship with color. Hyperspectral cameras can estimate SSC with high accuracy under controlled laboratory conditions, but their practicality in field environments is limited. This study investigates the potential of simple RGB sensors under uncontrolled lighting to estimate SSC and color, enabling cost-effective, robot-assisted harvesting. Over the 2021 and 2022 summer seasons, we collected grape images with corresponding SSC and color labels to evaluate algorithmic solutions for SSC estimation on embedded devices commonly used in robotics and smartphones. Our results demonstrate that SSC can be estimated from visual appearance with human-like performance. We propose computationally efficient histogram-based methods for resource-constrained robots and deep learning approaches for more complex applications.
@@
[inst1]organization=Sapienza University of Rome Department of Computer, Control and Management Engineering ,city=Rome, postcode=00185, state=RM, country=Italy
@@ 1 Introduction
The use of passive imaging sensors to measure agronomic properties of crops enables fast, non destructive monitoring that, when integrated with machinery or robots, can be efficiently scaled to cover entire fields. RGB cameras are commonly employed to monitor external conditions, such as detecting pests and diseases, while multi-spectral cameras are occasionally used to derive indices related to vegetation health. However, assessing internal crop properties presents a more complex challenge. Hyperspectral imagery (HSI) show potential in this area, as their ability to capture specific wavelength responses can be correlated with the absorption patterns of certain chemical properties. Nevertheless, hyperspectral cameras are typically slow to acquire images, highly sensitive to illumination conditions, and usually unsuitable for use on operational agricultural machinery.
Figure 1: Computer Vision plays a crucial role in enabling automatic harvesting operations; however, the requirements for high-quality performance often conflict with the economic and practical constraints of robot design. This work demonstrates the development and deployment of a simple RGB-based algorithm for SSC and color quality estimation, addressing the complexities of field operations while utilizing cost-effective sensors.
@@
These competing aspects collide with the desirable tension to bring more robots on the farming fields, in order to both face the shortage of workforce, but also to manage the crops in a more precise way. In fact, many cultivations are labor-intensive, and for various economic and social reasons, it is increasingly difficult to find human workers to perform operations such as harvesting. Table grape cultivation is an example of such labor-intensive cultivation that is increasingly popular in the markets but faces a workforce shortage. This shortage would be mitigated by the introduction of robot workers that could, to some extent, assist human workers in the field. To be able to do so, the robotic platform should be able to assess the quality and ripeness of the fruit using mainly passive sensors (active manipulation for perception would be orders of magnitude slower). While solutions to equip robots with hyperspectral cameras are becoming more prevalent, their high cost and operational complexity remain major barriers to widespread adoption. So, the question of designing a passive quality estimation system using low-cost, reliable, RGB cameras, or similar devices is still open.
@@
This work addresses this question by validating its feasibility and proposing high-performance algorithmic solutions tailored for practical deployment in agricultural settings. In the following paragraphs, we provide a brief review of the relevant literature, highlighting key contributions to contextualize this study. For a broader overview of fruit ripeness estimation methods—particularly for grapes using hyperspectral and RGB imaging—we refer the reader to [11, 17].
1.0.1 Hyperspectral applications to agricultural robotics
@@
Robots have been used for the monitoring and quality estimation of different crops and are now finding some industrial applications in high-value crops (wine grapes [20, 12], strawberries [23, 22], blueberries [19], among others). Some authors have explored specifically the quality estimation of fruits with hyperspectral sensors. In [6] an HSI device is mounted on an AGV to measure the TSS of Tempranillo wine grapes. The AGV moves along the vineyard row acquiring spectral data in a push-broom fashion. To achieve color constancy the authors perform an indirect calibration for exposure, while illuminant calibration was performed every 5m with a white reference. The impact of these choices on the accuracy of the estimation is not clear, and the authors report an RMS on °Brix estimate as low as $1.274\textdegree$. In addition, the method requires considerable post-processing time (5 hours), making it not well suited for robotic harvesting.
[18] propose a trailing cart mounting two HSI devices to monitor raspberries. However, the measurement is taken using an overarching arm from which a white reference is hung. [16] propose the Ladybug robot equipped with an HSI device built for low horticultural crops. The robot AGV is a straddle model, where the illumination constancy is granted through the use of a stroboscopic light.
@@
In preparation for robotic use, [9] perform a very thorough experimentation using VIS-SWIR images to estimate SSC in four wine grapes varieties. The authors analyze the contribution of different wavelengths to the estimation algorithms in depth and apply various algorithms for the prediction. Across the proposed models (Partial Last Square, Random Forest, Convolutional Neural Networks, and Support Vector Machines) and four grape varieties they obtain an RMS error ranging from 1.70 to 2.20 °Brix.
In a follow-up study, to address the problem of illuminant estimation, [15] propose to learn the illuminant normalization trough the use of an auto-encoder. The solution is similar to ours, in the sense that it learns the typical illuminant distribution, but loses the absolute metric accuracy of having a white reference in the scene. The authors explore four varieties of wine grapes, declaring RMSE values that range from 1.66 to 2.20 °Brix.
@@
1.0.2 RGB applications to fruit quality assessment
Given the limitations of HSI illuminant and color calibration and the cost of hyperspectral devices, many authors proposed approaches based on standard RGB sensors. [1] have explored the use of color imaging combined with advanced numerical techniques, such as interval numbers (INs), and neural network regressors to estimate grape maturity indices (e.g., TSS, TA, and pH), demonstrating the feasibility of non-destructive and cost-effective methods for integration into autonomous harvesting systems. In this work, the authors declare a considerable 6.6259 MSE Baumé, which approximately corresponds to an RMSE of 4.63 °Brix.
[21] perform a very interesting analytical study on 180 samples of Kyoho grape berries using an RGB camera. The authors propose the use of some feature indexes, similar in principle to the well-known NDVI, but computed with the standard RGB channels, obtaining a very low cross-validation error of 0.78 °Brix (RMSE) on a dataset with a standard deviation of 1.37°. The study is performed in laboratory conditions, but it is nonetheless significant for our work, as it shows the effectiveness of data-driven estimations with simple RGB cameras.
[3] proposed a non-destructive estimation technique based on an RGB camera using random forest predictors trained on color histograms. The histogram was computed in CIELab color space by clustering the colors and then choosing the ones with the highest intrinsic variance. The classification task was performed on a 5-level quality scale, and no explicit SSC measurement was included in the labeling process. The experimental setup involved 200 bunches, evenly divided between Italia and Victoria varieties, shot with artificial illumination, and image features were computed after background subtraction.
@@ 1.0.2 RGB applications to fruit quality assessment
Building on the challenges and opportunities highlighted in previous works, this study focuses on the problem of perception for quality estimation in table grape harvesting. Accurate quality estimation can enable robots to assist human workers or even surpass their capabilities during field operations. In high-quality table grape production, harvesting decisions typically follow a two-step process: an agronomist first samples grape bunches across the vineyard using a refractometer to measure SSC in °Brix. If the °Brix is sufficiently high, additional visual characteristics, such as color intensity and distribution, are evaluated. Workers are then instructed to harvest grape bunches based solely on appearance and location within the vineyard. While °Brix and color are not causally linked, they share latent ripening factors [8], allowing human experts to approximate °Brix from color to some extent. This work aims to replicate and enhance this human capability using data-driven methods and low-cost RGB sensors.
The contributions of our study are as follows:
@@
    A dataset of table grape images collected across different seasons and with different sensors, labeled with both °Brix values and color categories following the aforementioned agronomic procedure, thus enabling studies on SSC estimation in realistic field conditions.
@@
    Experiments and user studies to explore the extent to which °Brix can be estimated from color alone. These insights informed the design of algorithms and highlighted key factors for future research.
@@
    Two different estimation algorithms have been proposed and validated:
@@
      A lightweight method based on handcrafted features for constrained computational systems.
@@
      A deep, fully convolutional multi-task architecture that achieves high performance with limited training data.
  • 4.
    The methods do not require complex calibration and are designed for real-time use in robotic harvesting or mobile devices.
The paper is structured as follows: Section 2 describes the dataset that was collected and the estimation methods that were designed; Section 3 defines the experimental setup, and describes and comments on the experimental validation of the problem and the algorithms. Finally, Section 4 draws conclusions and outlines a path for future work.
@@ 2.1.1 Experimental Field
The experimental field is located in southern Lazio, Italy. The vineyard consists of two plots measuring approximately 114 m x 51 m (0.58 ha) and 122 m x 48 m (0.58 ha). The vineyards are structured as a traditional trellis system known as Tendone, with a wide distance between each plant (3 x 3 m). All structures are traditionally covered with plastic and netting to protect grapes from rain and hail. The plantations are all older than three years, healthy, and in full production, thus representing a typical working condition for the validation of agronomic activities such as fruit harvesting or vine pruning.
Among the varieties cultivated in this field, Black Pizzutello (Pizzutello nero), a distinct variety from the Lazio region, was selected due to its importance for the regional economy and color variability.

### v2

@@ Abstract
In table grape cultivation, harvesting depends on accurately assessing fruit quality. While some characteristics, like color, are visible, others, such as Soluble Solid Content (SSC), or sugar content measured in degrees Brix (°Brix), require specific tools. SSC is a key quality factor that correlates with ripeness, but lacks a direct causal relationship with color. Hyperspectral cameras can estimate SSC with high accuracy under controlled laboratory conditions, but their practicality in field environments is limited. This study investigates the potential of simple RGB sensors under uncontrolled lighting to estimate SSC and color, enabling cost-effective, robot-assisted harvesting. Over the 2021 and 2022 summer seasons, we collected grape images with corresponding SSC and color labels to evaluate algorithmic solutions for SSC estimation, specifically testing for cross-seasonal and cross-device robustness. We propose two approaches: a computationally efficient histogram-based method for resource-constrained robots and a Deep Neural Network (DNN) model for more complex applications. Our results demonstrate high performance, with the DNN model achieving a Mean Absolute Error (MAE) as low as $1.05$Brix on a challenging cross-device test set. The lightweight histogram-based method also proved effective, reaching an MAE of $1.46$Brix. These results are highly competitive with those from hyperspectral systems, which report errors in the 1.271.271.271.272.202.202.202.20 °Brix range in similar field applications.
@@
[…]
@@ 1 Introduction
The use of passive imaging sensors to measure agronomic properties of crops enables fast, non-destructive monitoring that, when integrated with machinery or robots, can be efficiently scaled to cover entire fields (Fig. 1). RGB (Red, Green, Blue) cameras are commonly employed to monitor external conditions, such as detecting pests and diseases, while multi-spectral cameras are occasionally used to derive indices related to vegetation health. However, assessing internal crop properties presents a more complex challenge. Hyperspectral imagery (HSI) shows potential in this area, as its ability to capture specific wavelength responses can be correlated with the absorption patterns of certain chemical properties. Nevertheless, hyperspectral cameras are typically slow to acquire images, highly sensitive to illumination conditions, and usually unsuitable for use on operational agricultural machinery.
Figure 1: A robotic platform for grape harvesting developed within the CANOPIES project ([1]), which provides the context for this research. Such robots require robust perception systems to assess fruit quality in the field using on-board sensors The end-effector and the head of the robot are equipped with simple, low-cost RGB cameras suitable for this task. This work focuses on developing and validating the core data-driven algorithms necessary for SSC and color estimation in these challenging robotic setups.
Table 1: Comparison of Hyperspectral (HSI) and RGB Imaging devices for in-field use.
Feature Hyperspectral Imaging (HSI) Standard RGB Imaging
Spectral Resolution High (hundreds of narrow bands, usually 5–10 nm) Low (3 broad bands: Red, Green, Blue)
Spatial Dimension Low High
Information Captured Detailed spectral signatures, correlated with chemical composition Surface color, texture, and shape
Equipment Cost $10000$$200000+$ $10$$1000+$
Operational Environment Laboratory or highly controlled settings In-field or indoors alike
Acquisition Time High (seconds to minutes, depending on illumination) Low (few milliseconds)
Computational Load High (due to the processing 3D hypercubes) Low (suitable for real-time processing on embedded systems)
Portability & Size Often bulky and requires careful calibration Compact, and easily integrated into robotic platforms
@@
These competing aspects are at odds with the desirable tension to bring more robots to the farming fields, in order to both face the shortage of workforce, but also to manage the crops in a more precise way. In fact, many cultivations are labor-intensive, and for various economic and social reasons, it is increasingly difficult to find human workers to perform operations such as harvesting. Table grape cultivation is an example of such labor-intensive cultivation that is increasingly popular in the markets but faces a workforce shortage. This shortage would be mitigated by the introduction of robot workers that could, to some extent, assist human workers in the field. To be able to do so, the robotic platform should be able to assess the quality and ripeness of the fruit, such as color, Soluble Solid Content (SSC), size, presence of anomalies, using mainly passive sensors (active manipulation for perception would be orders of magnitude slower). The case of SSC estimation is of particular relevance as its estimation via passive sensors is usually done with Infrared (IR) refractometers (that require at least contact with the berries) or hyperspectral imaging (HSI) devices. Although solutions equipping robots with HSI cameras are becoming more prevalent, their high cost and operational complexity remain major barriers to widespread adoption compared to the use of other devices, such as RGB cameras (Table 1 offers a summary of the main differences between hyperspectral and RGB cameras). So, the question of designing a passive SSC estimation system using low-cost, reliable, RGB cameras, or similar devices is still open.
@@
This work addresses this question by validating its feasibility and proposing high-performance algorithmic solutions tailored for practical deployment in agricultural settings. In the following paragraphs, we provide a brief review of the relevant literature, highlighting key contributions to contextualize this study. For a broader overview of fruit ripeness estimation methods—particularly for grapes using hyperspectral and RGB imaging—we refer the reader to Palumbo et al., [12] and Vrochidou et al., [20].
Table 2: Comparison of different approaches for SSC estimation in Table Grapes using spectral equipment.
Name Camera model Camera type Spectral Range Spectral Res. Spatial Res. Cultivar Illumination
Gutiérrez et al., [6] Resonon Pika Push broom VIS-NIR (400-1000nm) 2.1nm (300 bands) 300 px Tempranillo Sunlight
Kalopesa et al., [10] Contact probe spectrometer Spectrometer (contact probe) VNIR–SWIR (350–2500 nm) Not specified N/A (point measurement) Chardonnay, Malagouzia, Sauvignon-Blanc, Syrah Internal illumination
Tsakiridis et al., [18] Cubert FireflEYE V185 Snapshot VIS-NIR (400–1000 nm) 4-8nm (128 bands) 50x50 Chardonnay, Malagouzia, Sauvignon-Blanc, Syrah Sunlight with learned calibration
1.1 Hyperspectral applications to agricultural robotics
Robots have been used for the monitoring and quality estimation of different crops and are now finding some industrial applications in high-value crops (e.g. wine grapes in Williams et al., [23] and Polvara et al., [13], strawberries in Xiong et al., [26, 25], blueberries in Williams et al., [22], among others). Some authors have explored specifically the quality estimation of fruits with hyperspectral sensors. In Table 2 we summarize the HSI approaches applied to grapes internal quality estimation.
@@ 1.1 Hyperspectral applications to agricultural robotics
In Gutiérrez et al., [6] an HSI device is mounted on an Autonomous Ground Vehicle (AGV) to measure the Total Soluble Solids (TSS) — a synonym of SSC — of Tempranillo wine grapes. The AGV moves along the vineyard row acquiring spectral data in a push-broom fashion. To achieve color constancy the authors perform an indirect calibration for exposure, while illuminant calibration was performed every 5m with a white reference. Although the authors claim that this setup shows on-the-go HSI estimation feasibility, they also note the difficulty of maintaining the steady motion required by push-broom cameras using an AGV. In addition, white balancing stop-and-go operations are required every 5 meters to calibrate the camera. The impact of these choices on the accuracy of the estimation is not clear, and the authors report a Root Mean Square Error (RMSE) on °Brix estimate as low as $1.274$°. Finally, the method requires considerable post-processing time (5 hours), making it not well suited for robotic harvesting. Being able to use a simple RGB camera for the estimation would remove both the vibration and calibration issues.
A similar approach, but with a trailing cart instead of an AGV, was proposed by Williams et al., [21]. The cart mounts two HSI devices (VNIR: 400-896nm, SWIR: 896-2506nm) to monitor raspberries. Again, the push broom image is obtained by moving a vehicle and the whole setup requires an overarching arm with a black background screen behind the plants and from which a white reference is hung. The aim of the study was to just demonstrate the setup and no physical characteristics have been estimated using the hyperspectral data.
A very successful robotic platform, called Ladybug, is presented by Underwood et al., [19]. The Ladybug robot, among other sensors and navigation equipment, is also equipped with an HSI device (Resonon Pika II, VNIR: 390-887nm) built for low horticultural crops. The robot AGV is a straddle model, where the illumination constancy is granted through the use of a stroboscopic light, in order to do not interfere with the rest of the optical sensors. In addition, calibration pads have been placed both on the field and on the robot, on the periphery of the HSI device’s field of view (FOV). The camera has been used to compute only the Normalized Difference Vegetation Index (NDVI), as the whole application was built to gather phenotyping information such as the plant height. The use of straddle systems is one of the most promising strategies for the application of HSI in the fields, however the calibration challenges remain, and the use for taller crops is unpractical.
@@
In preparation for robotic use, Kalopesa et al., [10] perform a very thorough experimentation using VIS-SWIR contact spectrometer (400-2500nm) to estimate SSC in four wine grapes varieties, namely Chardonnay, Malagouzia, Sauvignon-Blanc and Syrah. Contact probe spectrometers have a similar output to HSI imaging devices, but are meant for precise point sampling. These devices provide an internal illumination source and have to be carried around with a handle, or in a backpack. The cost is still higher than RGB cameras, and more complex to integrate into a robotic platform due to the delicate physical contact requirement. The authors analyze the contribution of different wavelengths to the estimation algorithms in depth and apply various algorithms for the prediction. Across the proposed models – Partial Least Square Regression (PLSR), Random Forest (RF), Convolutional Neural Networks (CNN), and Support Vector Machines (SVM) – and four grape varieties they obtain an RMSE ranging from 1.70 to 2.20 °Brix.
In a follow-up study (Tsakiridis et al., [18]), in addition to the spectrometer, the authors introduce the in-field estimation of SSC with the use of a snapshot HSI camera (the Cubert FireflEYE V185, VNIR: 400-1000nm). In a different fashion from other works the main contribution of this work is a method to address the problem of illuminant estimation without the use of a reference tile before each acquisition. Instead they propose to learn the illuminant normalization through the use of an auto-encoder. The ground truth for the SSC estimation is collected with an analog refractometer, sampling the °Brix of a single berry for each bunch. The solution of using a deep neural network to estimate the illuminant distribution bears some similarities with our work as it simplifies the acquisition setup, but loses the absolute metric accuracy of having a white reference in the scene and still requires periodic re-calibration of the network using an more reliable spectrometer. In addition, the use of a snapshot HSI camera considerably reduces the spatial resolution, with respect to push broom setups. The authors explore four varieties of wine grapes, declaring RMSE values that range from 1.66 to 2.20 °Brix.
@@
1.2 RGB applications to fruit quality assessment
Given the limitations of HSI illuminant and color calibration, and the cost of hyperspectral devices, many authors proposed approaches based on standard RGB sensors. Bazinas et al., [2] have explored the use of color imaging combined with advanced numerical techniques, such as interval numbers (INs), and neural network regressors to estimate grape maturity indices (e.g., Total Soluble Solids (TSS), Titratable Acidity (TA), and pH), demonstrating the feasibility of non-destructive and cost-effective methods for integration into autonomous harvesting systems. The authors use a stereo camera (ZED Mini) to acquire RGB images converted to CIELab color space, under natural daylight. The grape variety is Tempranillo that, as in our case, is a black grape cultivar, that allows for an easier correlation of color and SSC. While similar to our approach, the declared performances are considerably higher, i.e. 6.6259 Baumé Mean Squared Error (MSE), which approximately corresponds to an RMSE of 4.63 °Brix.
Xia et al., [24] perform a very interesting analytical study on 180 samples of Kyoho grape berries using Samsung industrial RGB camera (SCC-B1011). The authors propose the use of some feature indices, similar in principle to the well-known NDVI, but computed with the standard RGB channels, as the inputs to two estimation methods, Partial Least Squares Regression (PLSR) and Multiple Linear Regression (MLR), obtaining a very low cross-validation error of 0.78 °Brix (RMSE) on a dataset with a standard deviation of 1.37°. The study is performed under laboratory conditions, but it is nonetheless significant for our work, as it shows the effectiveness of data-driven estimations with simple RGB cameras. In this work we show that also in uncontrolled environments it is possible to achieve estimation errors low enough to perform automatic harvesting reliably.
Table 3: Summary of RGB-based approaches for fruit quality assessment (focus on SSC).
Reference Camera Model Cultivar(s) Illumination Key Method / Features SSC Performance
Bazinas et al., [2] ZED Mini 3D Tempranillo (100 samples of 10 vine trees in multiple dates) Daylight Color imaging, Interval Numbers (INs), Neural Network regressors RMSE 4.63 °Brix
Xia et al., [24] Samsung SCC-B1011 camera Kyoho (180 berry samples) Laboratory conditions Custom RGB feature indices; PLSR and MLR RMSE 0.78 °Brix (Cross-Validation).
Cavallo et al., [4] RGB camera Italia, Victoria (200 bunches) Artificial illumination CIELab color histograms (clustered, high variance colors), Random Forest, background subtraction. 5-level quality scale.
Cavallo et al., [4] proposed a non-destructive estimation technique based on an RGB camera (JAI CV-M9GE industrial CCD camera) using random forest predictors trained on color histograms. The histogram was computed in CIELab color space by clustering the colors and then choosing the ones with the highest intrinsic variance. While this is the only study that targets white grapes, the estimation problem was framed as a qualitative color scale measurement. The classification task was performed on a 5-level quality scale, and no explicit SSC measurement was included in the labeling process. However, the harvested bunches provided were nominally above 16° Brix level, that for the table grape food industry is a common quality requirement. The experimental setup involved 400 bunches, evenly divided between Italia and Victoria varieties, shot with artificial illumination, and image features were computed after background subtraction and careful white balancing using color calibration references. While this study cannot be quantitatively compared to ours, it is noteworthy as it shows that even with white grape varieties the correlation between color and SSC can be captured by careful modeling. The main limitation of this work that we are trying to overcome is the work in a controlled environment. For this reason, in this initial study we focus our attention to a black grape variety, as in Bazinas et al., [2] and Xia et al., [24]. Table 3 summarizes these RGB-based approaches to SSC estimation in grapes.
@@ 1.2 RGB applications to fruit quality assessment
This study addresses the challenge of perceptual quality estimation for robotic table grape harvesting, a task that could enable robots to assist or even surpass human workers. We model our approach on the standard harvesting procedure where an agronomist first confirms a target SSC in °Brix using a refractometer. Subsequently, workers harvest individual bunches based on visual cues like color. Although SSC and color are not causally linked, they share latent ripening factors – such as climate, weather, training practices and use of nitrogen, to name a few (Jackson, [9]) – which allows human experts to visually approximate ripeness. Our central goal is to replicate and enhance this expert capability using data-driven methods with low-cost RGB sensors.
Our primary contributions are:
@@
    A multi-season, multi-sensor dataset of in-field grape images, labeled with both SSC (°Brix) and agronomic color categories. To the best of our knowledge, this is the first publicly available dataset that combines images with °Brix labeling, thus enabling studies on quality estimation under realistic field conditions.
@@
    An analysis of the visual-SSC correlation, including a user study that establishes a baseline for human-level performance in estimating ripeness from images.
@@
    The design and validation of two distinct algorithms for real-time, calibration-free operation:
@@
      A lightweight, histogram-based method for resource-constrained robotic systems.
@@
      A deep, fully convolutional network that achieves high performance even with limited data.
The remainder of this paper is structured as follows. Section 2 details our dataset and proposed estimation methods. Section 3 presents the experimental validation and and Section 4 discusses performance, limitations and computational constraints. Finally, Section 5 concludes the paper and outlines future research directions.
@@ 2 Materials and Methods
In this Section we detail the data acquisition and labeling processes and the two algorithmic approaches to SSC estimation. Fig, 2 shows the general overview of the methodology.
Figure 2: An overview of the experimental workflow presented in this paper. The process moves from left to right: (1) In-field Data Collection using multiple RGB sensors and subsequent labeling for SSC and color quality. (2) Organization of the data into three distinct Datasets to evaluate performance under different conditions. (3) Development and validation of two estimation algorithms: a lightweight histogram-based model and a high-performance multi-task DNN. (4) The Test panel summarizes the key result, outlining our best model’s SSC estimation error (MAE) and harvesting decision performance (Precision/Recall) against the human expert baseline.
@@ 2.1.1 Experimental Field
The experimental field is located in southern Lazio, Italy. The vineyard consists of two plots measuring approximately 114 m x 51 m (0.58 ha) and 122 m x 48 m (0.58 ha). The vineyards are structured as a traditional trellis system known as Tendone, with a wide distance between each plant (3 x 3 m). All structures are traditionally covered with plastic and netting to protect grapes from rain and hail. The plantations are all older than three years, healthy, and in full production, thus representing a typical working condition for the validation of agronomic activities such as fruit harvesting or vine pruning. Among the varieties cultivated in this field, Black Pizzutello (Pizzutello nero), a distinct variety from the Lazio region, was selected due to its importance for the regional economy and color variability.